OpenStudy (anonymous):

simplify the radical expression sqrt20

OpenStudy (hba):

How would you go about doing this?

OpenStudy (anonymous):

im not sure

OpenStudy (solomonzelman):

\[Hint:~~20=2^2 \times 5\]

OpenStudy (anonymous):

so squareroot 2x 5?

OpenStudy (solomonzelman):

I am going to give you an example. \[\sqrt{48}=\sqrt{16 \times 3}=\sqrt{4^2 \times 3}=\sqrt{4^2} \times \sqrt{3}=4\sqrt{3}\]So, \[\sqrt{20}=\sqrt{4 \times 5}=\sqrt{2^2 \times 5}=??\]

OpenStudy (solomonzelman):

Can you finish the problem based on the example ?

OpenStudy (anonymous):

2sqareroot5?

OpenStudy (solomonzelman):

Yes.. that's right !!

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

do you think you could help with another one?